# Repeating Flares, X-ray Outbursts and Delayed Infrared Emission: A Comprehensive Compilation of Optical Tidal Disruption Events

### v1

@@ Acknowledgements.
- Fruscione et al. (2006) Fruscione, A., McDowell, J. C., Allen, G. E., et al. 2006, in Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 6270, Observatory Operations: Strategies, Processes, and Systems, ed. D. R. Silva & R. E. Doxsey, 62701V
- Gehrels (1986) Gehrels, N. 1986, ApJ, 303, 336
- Gezari et al. (2012) Gezari, S., Chornock, R., Rest, A., et al. 2012, Nature, 485, 217
- Gilfanov et al. (2020) Gilfanov, M., Sazonov, S., Sunyaev, R., et al. 2020, The Astronomer’s Telegram, 14246, 1
- Götberg et al. (2019) Götberg, Y., de Mink, S. E., Groh, J. H., Leitherer, C., & Norman, C. 2019, A&A, 629, A134
- Graham et al. (2019) Graham, M. J., Kulkarni, S. R., Bellm, E. C., et al. 2019, PASP, 131, 078001
- Grupe et al. (1999) Grupe, D., Thomas, H. C., & Leighly, K. M. 1999, A&A, 350, L31
- Guo et al. (2025) Guo, H., Sun, J., Li, S., et al. 2025, ApJ, 979, 235
- Hammerstein et al. (2023) Hammerstein, E., van Velzen, S., Gezari, S., et al. 2023, ApJ, 942, 9
- Harris et al. (2020) Harris, C. R., Millman, K. J., Van Der Walt, S. J., et al. 2020, Nature, 585, 357
- Hart et al. (2023) Hart, K., Shappee, B. J., Hey, D., et al. 2023, arXiv e-prints, arXiv:2304.03791
- Heinze et al. (2018) Heinze, A. N., Tonry, J. L., Denneau, L., et al. 2018, AJ, 156, 241
- Hills (1975) Hills, J. G. 1975, Nature, 254, 295
- Hinkle et al. (2021) Hinkle, J. T., Holoien, T. W. S., Auchettl, K., et al. 2021, MNRAS, 500, 1673
- Ho et al. (2025) Ho, A. Y. Q., Yao, Y., Matsumoto, T., et al. 2025, arXiv e-prints, arXiv:2502.07885
- Hodgkin et al. (2021) Hodgkin, S. T., Harrison, D. L., Breedt, E., et al. 2021, A&A, 652, A76
- Holoien et al. (2016a) Holoien, T. W. S., Kochanek, C. S., Prieto, J. L., et al. 2016a, MNRAS, 463, 3813
- Holoien et al. (2016b) Holoien, T. W. S., Kochanek, C. S., Prieto, J. L., et al. 2016b, MNRAS, 455, 2918
- Holoien et al. (2014) Holoien, T. W. S., Prieto, J. L., Bersier, D., et al. 2014, MNRAS, 445, 3263
- Hung et al. (2017) Hung, T., Gezari, S., Blagorodnova, N., et al. 2017, ApJ, 842, 29
- Hunter (2007) Hunter, J. D. 2007, Computing in Science and Engineering, 9, 90
- Iben & Tutukov (1985) Iben, Jr., I. & Tutukov, A. V. 1985, ApJS, 58, 661
- Inserra et al. (2012) Inserra, C., Smartt, S. J., Fraser, M., et al. 2012, The Astronomer’s Telegram, 4329, 1
- Ivezic et al. (2008) Ivezic, Z., Axelrod, T., Brandt, W. N., et al. 2008, Serbian Astronomical Journal, 176, 1
- Jiang et al. (2016) Jiang, N., Dou, L., Wang, T., et al. 2016, ApJ, 828, L14
- Jiang et al. (2021) Jiang, N., Wang, T., Dou, L., et al. 2021, ApJS, 252, 32
- Jose et al. (2014) Jose, J., Guo, Z., Long, F., et al. 2014, The Astronomer’s Telegram, 6777, 1
- Kalberla et al. (2005) Kalberla, P. M. W., Burton, W. B., Hartmann, D., et al. 2005, A&A, 440, 775
- Khorunzhev et al. (2022) Khorunzhev, G. A., Sazonov, S. Y., Medvedev, P. S., et al. 2022, Astronomy Letters, 48, 767
- Kochanek (2016) Kochanek, C. S. 2016, MNRAS, 461, 371
- Kochanek et al. (2017) Kochanek, C. S., Shappee, B. J., Stanek, K. Z., et al. 2017, PASP, 129, 104502
- Koljonen et al. (2024) Koljonen, K. I. I., Liodakis, I., Lindfors, E., et al. 2024, MNRAS, 532, 112
- Krimm et al. (2011) Krimm, H. A., Kennea, J. A., Holland, S. T., et al. 2011, The Astronomer’s Telegram, 3384, 1
- Kumar et al. (2024) Kumar, H., Berger, E., Hiramatsu, D., et al. 2024, ApJ, 974, L36
- Law et al. (2009) Law, N. M., Kulkarni, S. R., Dekany, R. G., et al. 2009, PASP, 121, 1395
- Li et al. (2024) Li, R.-L., Yuan, C., He, H.-N., et al. 2024, arXiv e-prints, arXiv:2411.06440
- Liddle (2007) Liddle, A. R. 2007, MNRAS, 377, L74
- Lin et al. (2024) Lin, Z., Jiang, N., Wang, T., et al. 2024, ApJ, 971, L26
- Liu et al. (2025) Liu, C., Yarza, R., & Ramirez-Ruiz, E. 2025, ApJ, 979, 40
- Liu et al. (2017) Liu, T., Tozzi, P., Wang, J.-X., et al. 2017, ApJS, 232, 8
- Magnier et al. (2020a) Magnier, E. A., Chambers, K. C., Flewelling, H. A., et al. 2020a, ApJS, 251, 3
- Magnier et al. (2020b) Magnier, E. A., Schlafly, E. F., Finkbeiner, D. P., et al. 2020b, ApJS, 251, 6
- Magnier et al. (2020c) Magnier, E. A., Sweeney, W. E., Chambers, K. C., et al. 2020c, ApJS, 251, 5
- Mainetti et al. (2016) Mainetti, D., Lupi, A., Campana, S., & Colpi, M. 2016, MNRAS, 457, 2516
- Mainzer et al. (2014) Mainzer, A., Bauer, J., Cutri, R. M., et al. 2014, ApJ, 792, 30
- Mainzer et al. (2011) Mainzer, A., Bauer, J., Grav, T., et al. 2011, ApJ, 731, 53
- Makrygianni et al. (2025) Makrygianni, L., Arcavi, I., Newsome, M., et al. 2025, arXiv e-prints, arXiv:2505.16867
- Mandel & Levin (2015) Mandel, I. & Levin, Y. 2015, ApJ, 805, L4
- Marchesi et al. (2016) Marchesi, S., Lanzuisi, G., Civano, F., et al. 2016, ApJ, 830, 100
- Masci et al. (2019) Masci, F. J., Laher, R. R., Rusholme, B., et al. 2019, PASP, 131, 018003
- Masterson et al. (2024) Masterson, M., De, K., Panagiotou, C., et al. 2024, ApJ, 961, 211
- Merloni et al. (2024) Merloni, A., Lamer, G., Liu, T., et al. 2024, A&A, 682, A34
- Miller et al. (2015) Miller, J. M., Kaastra, J. S., Miller, M. C., et al. 2015, Nature, 526, 542
- Monet et al. (2003) Monet, D. G., Levine, S. E., Canzian, B., et al. 2003, AJ, 125, 984
@@
- Moretti et al. (2005) Moretti, A., Campana, S., Mineo, T., et al. 2005, in Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 5898, UV, X-Ray, and Gamma-Ray Space Instrumentation for Astronomy XIV, ed. O. H. W. Siegmund, 360–368
- Munoz-Arancibia et al. (2023) Munoz-Arancibia, A., Hernandez-Garcia, L., Bauer, F. E., et al. 2023, Transient Name Server AstroNote, 138, 1
- Nevalainen et al. (2005) Nevalainen, J., Markevitch, M., & Lumb, D. 2005, ApJ, 629, 172
- Newsome et al. (2022) Newsome, M., Arcavi, I., Dgany, Y., & Pellegrino, C. 2022, Transient Name Server AstroNote, 236, 1
- Newsome et al. (2024) Newsome, M., Arcavi, I., Howell, D. A., et al. 2024, ApJ, 977, 258
- Nicholl et al. (2019) Nicholl, M., Blanchard, P. K., Berger, E., et al. 2019, MNRAS, 488, 1878
- Nicholl et al. (2022) Nicholl, M., Lanning, D., Ramsden, P., et al. 2022, MNRAS, 515, 5604
- Oke & Gunn (1983) Oke, J. B. & Gunn, J. E. 1983, ApJ, 266, 713
- Onori et al. (2022) Onori, F., Cannizzaro, G., Jonker, P. G., et al. 2022, MNRAS, 517, 76
- Paczyński (1967) Paczyński, B. 1967, Acta Astron., 17, 355
- Paggi et al. (2013) Paggi, A., Massaro, F., D’Abrusco, R., et al. 2013, ApJS, 209, 9
- Panagiotou et al. (2023) Panagiotou, C., De, K., Masterson, M., et al. 2023, ApJ, 948, L5
- Pasham et al. (2017) Pasham, D. R., Cenko, S. B., Sadowski, A., et al. 2017, ApJ, 837, L30
- Pasham et al. (2023) Pasham, D. R., Lucchini, M., Laskar, T., et al. 2023, Nature Astronomy, 7, 88
- Perri et al. (2007) Perri, M., Maselli, A., Giommi, P., et al. 2007, A&A, 462, 889
- Poidevin et al. (2023) Poidevin, F., Pérez-Fournon, I., Geier, S., et al. 2023, Transient Name Server Classification Report, 2023-3050, 1
- Prieto et al. (2014) Prieto, J. L., Bersier, D., Holoien, T. W. S., et al. 2014, The Astronomer’s Telegram, 5831, 1
- Puccetti et al. (2011) Puccetti, S., Capalbi, M., Giommi, P., et al. 2011, A&A, 528, A122
- Rau et al. (2009) Rau, A., Kulkarni, S. R., Law, N. M., et al. 2009, PASP, 121, 1334
- Rees (1988) Rees, M. J. 1988, Nature, 333, 523
- Reis et al. (2012) Reis, R. C., Reynolds, M. T., Miller, J. M., Miller, B. P., & Hadjiyska, E. 2012, The Astronomer’s Telegram, 4346, 1
- Ricci et al. (2017) Ricci, C., Trakhtenbrot, B., Koss, M. J., et al. 2017, ApJS, 233, 17
- Roming et al. (2005) Roming, P. W. A., Kennedy, T. E., Mason, K. O., et al. 2005, Space Sci. Rev., 120, 95
- Saxton et al. (2020) Saxton, R., Komossa, S., Auchettl, K., & Jonker, P. G. 2020, Space Sci. Rev., 216, 85
- Sazonov et al. (2021) Sazonov, S., Gilfanov, M., Medvedev, P., et al. 2021, MNRAS, 508, 3820
- Scargle (1998) Scargle, J. D. 1998, ApJ, 504, 405
- Scargle et al. (2013) Scargle, J. D., Norris, J. P., Jackson, B., & Chiang, J. 2013, ApJ, 764, 167
- Schlafly & Finkbeiner (2011) Schlafly, E. F. & Finkbeiner, D. P. 2011, ApJ, 737, 103
- Shappee et al. (2014) Shappee, B. J., Prieto, J. L., Grupe, D., et al. 2014, ApJ, 788, 48
- Shingles et al. (2021) Shingles, L., Smith, K. W., Young, D. R., et al. 2021, Transient Name Server AstroNote, 7, 1
- Short et al. (2023) Short, P., Lawrence, A., Nicholl, M., et al. 2023, MNRAS, 525, 1568
- Smartt et al. (2012) Smartt, S. J., Inserra, C., Fraser, M., et al. 2012, The Astronomer’s Telegram, 4299, 1
- Smith et al. (2020) Smith, K. W., Smartt, S. J., Young, D. R., et al. 2020, PASP, 132, 085002
- Somalwar et al. (2023a) Somalwar, J. J., Ravi, V., Dong, D. Z., et al. 2023a, arXiv e-prints, arXiv:2310.03791
- Somalwar et al. (2023b) Somalwar, J. J., Ravi, V., Yao, Y., et al. 2023b, arXiv e-prints, arXiv:2310.03782
- Stein et al. (2021a) Stein, R., van Velzen, S., Kowalski, M., et al. 2021a, Nature Astronomy, 5, 510
- Stein et al. (2021b) Stein, R., van Velzen, S., Kowalski, M., et al. 2021b, Nature Astronomy, 5, 510
- Stokes et al. (2000) Stokes, G. H., Evans, J. B., Viggh, H. E. M., Shelly, F. C., & Pearce, E. C. 2000, Icarus, 148, 21
- Subrayan et al. (2023) Subrayan, B. M., Milisavljevic, D., Chornock, R., et al. 2023, ApJ, 948, L19

### v2

@@ Acknowledgements.
- CIAO: Chandra’s data analysis system. In Observatory Operations: Strategies, Processes, and Systems, D. R. Silva and R. E. Doxsey (Eds.), Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 6270, pp. 62701V. External Links: Document, ADS entry Cited by: §B.2.1.
- Confidence Limits for Small Numbers of Events in Astrophysical Data. ApJ 303, pp. 336. External Links: Document, ADS entry Cited by: §B.4.
- An ultraviolet-optical flare from the tidal disruption of a helium-rich stellar core. Nature 485 (7397), pp. 217–220. External Links: Document, 1205.0252, ADS entry Cited by: Appendix A, 1st item, 8th item, §5.2.
- Tidal disruption event SRGeJ213527.3-181634/ZTF20abgbdpr. The Astronomer’s Telegram 14246, pp. 1. External Links: ADS entry Cited by: Appendix A.
- The impact of stars stripped in binaries on the integrated spectra of stellar populations. A&A 629, pp. A134. External Links: Document, 1908.06102, ADS entry Cited by: §5.2.
- The Zwicky Transient Facility: Science Objectives. PASP 131 (1001), pp. 078001. External Links: Document, 1902.01945, ADS entry Cited by: §1.
- A Dependence of the Tidal Disruption Event Rate on Global Stellar Surface Mass Density and Stellar Velocity Dispersion. ApJ 853 (1), pp. 39. External Links: Document, 1707.02986, ADS entry Cited by: §4.3.
- The population of tidal disruption events discovered with eROSITA. A&A 697, pp. A159. External Links: Document, 2504.08424, ADS entry Cited by: §1.
- RX J1624.9+7554: a new X-ray transient AGN. A&A 350, pp. L31–L34. External Links: Document, astro-ph/9909101, ADS entry Cited by: §1.
- Reverberation Evidence for Stream Collision and Delayed Disk Formation in Tidal Disruption Events. ApJ 979 (2), pp. 235. External Links: Document, 2312.06771, ADS entry Cited by: §4.3, §5.3, footnote 16.
- The Final Season Reimagined: 30 Tidal Disruption Events from the ZTF-I Survey. ApJ 942 (1), pp. 9. External Links: Document, 2203.01461, ADS entry Cited by: §1, §2.1, §2.1, §2.1, Table 1, §3.4, §4.2, §4.2, §5.1, §5.2, §5.3, footnote 16.
- Array programming with numpy. Nature 585 (7825), pp. 357–362. Cited by: Data availability.
- ASAS-SN Sky Patrol V2.0. arXiv e-prints, pp. arXiv:2304.03791. External Links: Document, 2304.03791, ADS entry Cited by: §1.
- A First Catalog of Variable Stars Measured by the Asteroid Terrestrial-impact Last Alert System (ATLAS). AJ 156 (5), pp. 241. External Links: Document, 1804.02132, ADS entry Cited by: §1.
- Possible power source of Seyfert galaxies and QSOs. Nature 254 (5498), pp. 295–298. External Links: Document, ADS entry Cited by: §1.
- Hyper-velocity and tidal stars from binaries disrupted by a massive Galactic black hole. Nature 331 (6158), pp. 687–689. External Links: Document, ADS entry Cited by: §4.3.
- Discovery and follow-up of ASASSN-19dj: an X-ray and UV luminous TDE in an extreme post-starburst galaxy. MNRAS 500 (2), pp. 1673–1696. External Links: Document, 2006.06690, ADS entry Cited by: §5.3, footnote 16.
- A Luminous Red Optical Flare and Hard X-Ray Emission in the Tidal Disruption Event AT 2024kmq. ApJ 989 (1), pp. 54. External Links: Document, 2502.07885, ADS entry Cited by: Table 1.
- Gaia Early Data Release 3. Gaia photometric science alerts. A&A 652, pp. A76. External Links: Document, 2106.01394, ADS entry Cited by: §1.
- ASASSN-15oi: a rapidly evolving, luminous tidal disruption event at 216 Mpc. MNRAS 463 (4), pp. 3813–3828. External Links: Document, 1602.01088, ADS entry Cited by: Appendix A, 2nd item, 13rd item.
- Six months of multiwavelength follow-up of the tidal disruption candidate ASASSN-14li and implied TDE rates from ASAS-SN. MNRAS 455 (3), pp. 2918–2935. External Links: Document, 1507.01598, ADS entry Cited by: Appendix A, §1, 11st item.
- ASASSN-14ae: a tidal disruption event at 200 Mpc. MNRAS 445 (3), pp. 3263–3277. External Links: Document, 1405.1417, ADS entry Cited by: Appendix A, §1, 12nd item.
- Revisiting Optical Tidal Disruption Events with iPTF16axa. ApJ 842 (1), pp. 29. External Links: Document, 1703.01299, ADS entry Cited by: Appendix A, §1, 4th item, 10th item.
- Matplotlib: A 2D Graphics Environment. Computing in Science and Engineering 9 (3), pp. 90–95. External Links: Document, ADS entry Cited by: Data availability.
- On the evolution of close binaries with components of initial mass between 3 M and 12 M.. ApJS 58, pp. 661–710. External Links: Document, ADS entry Cited by: §5.2.
- Further spectral classification of PESSTO blue transients. The Astronomer’s Telegram 4329, pp. 1. External Links: ADS entry Cited by: Appendix A.
- Large Synoptic Survey Telescope: From Science Drivers To Reference Design. Serbian Astronomical Journal 176, pp. 1–13. External Links: Document, ADS entry Cited by: §5.1.
- The WISE Detection of an Infrared Echo in Tidal Disruption Event ASASSN-14li. ApJ 828 (1), pp. L14. External Links: Document, 1605.04640, ADS entry Cited by: Appendix A, §1.
- Mid-infrared Outbursts in Nearby Galaxies (MIRONG). I. Sample Selection and Characterization. ApJS 252 (2), pp. 32. External Links: Document, 2012.06806, ADS entry Cited by: §6.
- ASAS-SN Discovery of an Unusual Nuclear Transient in PGC 043234. The Astronomer’s Telegram 6777, pp. 1. External Links: ADS entry Cited by: Appendix A.
- The Leiden/Argentine/Bonn (LAB) Survey of Galactic HI. Final data release of the combined LDS and IAR surveys with improved stray-radiation corrections. A&A 440 (2), pp. 775–782. External Links: Document, astro-ph/0504140, ADS entry Cited by: §B.1.2.
- Search for Tidal Disruption Events Based on the SRG/eROSITA Survey with Subsequent Optical Spectroscopy. Astronomy Letters 48 (12), pp. 767–789. External Links: Document, ADS entry Cited by: §6.
- The All-Sky Automated Survey for Supernovae (ASAS-SN) Light Curve Server v1.0. PASP 129 (980), pp. 104502. External Links: Document, 1706.07060, ADS entry Cited by: §1.
- Tidal disruption event demographics. MNRAS 461 (1), pp. 371–384. External Links: Document, 1601.06787, ADS entry Cited by: §5.2.
- The extreme coronal line emitter AT 2022fpx: varying optical polarization properties and late-time X-ray flare. MNRAS 532 (1), pp. 112–125. External Links: Document, 2403.04877, ADS entry Cited by: §5.4.
- Swift reports the detection of a new transient source Swift J2058.4+0516. The Astronomer’s Telegram 3384, pp. 1. External Links: ADS entry Cited by: Appendix A.
- AT2023vto: An Exceptionally Luminous Helium Tidal Disruption Event from a Massive Star. ApJ 974 (2), pp. L36. External Links: Document, 2408.01482, ADS entry Cited by: Appendix A, 1st item.
- The Palomar Transient Factory: System Overview, Performance, and First Results. PASP 121 (886), pp. 1395. External Links: Document, 0906.5350, ADS entry Cited by: §1.
- Tidal Disruption Event Host Galaxies in the Context of the Local Galaxy Population. ApJ 850 (1), pp. 22. External Links: Document, 1707.01559, ADS entry Cited by: §4.3.
- A neutrino flare associated with X-ray emission from TDE ATLAS17jrp. arXiv e-prints, pp. arXiv:2411.06440. External Links: Document, 2411.06440, ADS entry Cited by: Appendix A, §5.4.
- Information criteria for astrophysical model selection. MNRAS 377 (1), pp. L74–L78. External Links: Document, astro-ph/0701113, ADS entry Cited by: §B.4.
- Discovery of an Ultrasoft X-Ray Transient Source in the 2XMM Catalog: A Tidal Disruption Event Candidate. ApJ 738 (1), pp. 52. External Links: Document, 1106.0744, ADS entry Cited by: §1.
- The Unluckiest Star: A Spectroscopically Confirmed Repeated Partial Tidal Disruption Event AT 2022dbl. ApJ 971 (1), pp. L26. External Links: Document, 2405.10895, ADS entry Cited by: footnote 16.
- Repeating Partial Tidal Encounters of Sun-like Stars Leading to Their Complete Disruption. ApJ 979 (1), pp. 40. External Links: Document, 2406.01670, ADS entry Cited by: §5.3.
- The Pan-STARRS Data-processing System. ApJS 251 (1), pp. 3. External Links: Document, 1612.05240, ADS entry Cited by: §1.
- Pan-STARRS Pixel Analysis: Source Detection and Characterization. ApJS 251 (1), pp. 5. External Links: Document, 1612.05244, ADS entry Cited by: §1.
- Pan-STARRS Photometric and Astrometric Calibration. ApJS 251 (1), pp. 6. External Links: Document, 1612.05242, ADS entry Cited by: §1.
- Rates of tidal disruption of stars by massive central black holes. MNRAS 309 (2), pp. 447–460. External Links: Document, astro-ph/9902032, ADS entry Cited by: §1.
- Hydrodynamical simulations of the tidal stripping of binary stars by massive black holes. MNRAS 457 (3), pp. 2516–2529. External Links: Document, 1601.05478, ADS entry Cited by: §4.3.
- Initial Performance of the NEOWISE Reactivation Mission. ApJ 792 (1), pp. 30. External Links: Document, 1406.6025, ADS entry Cited by: §1.
- Preliminary Results from NEOWISE: An Enhancement to the Wide-field Infrared Survey Explorer for Solar System Science. ApJ 731 (1), pp. 53. External Links: Document, 1102.1996, ADS entry Cited by: §1.
- The Double Tidal Disruption Event AT 2022dbl Implies that at Least Some “Standard” Optical Tidal Disruption Events Are Partial Disruptions. ApJ 987 (1), pp. L20. External Links: Document, 2505.16867, ADS entry Cited by: §5.3, §5.3, footnote 16.
- Double Tidal Disruptions in Galactic Nuclei. ApJ 805 (1), pp. L4. External Links: Document, 1504.02787, ADS entry Cited by: §4.3.
- The Zwicky Transient Facility: Data Processing, Products, and Archive. PASP 131 (995), pp. 018003. External Links: Document, 1902.01872, ADS entry Cited by: §1.
- A New Population of Mid-infrared-selected Tidal Disruption Events: Implications for Tidal Disruption Event Rates and Host Galaxy Properties. ApJ 961 (2), pp. 211. External Links: Document, 2401.01403, ADS entry Cited by: §2.2, §5.4, §6.
- The SRG/eROSITA all-sky survey. First X-ray catalogues and data release of the western Galactic hemisphere. A&A 682, pp. A34. External Links: Document, 2401.17274, ADS entry Cited by: §B.5.2, §3.2.
- Flows of X-ray gas reveal the disruption of a star by a massive black hole. Nature 526 (7574), pp. 542–545. External Links: Document, 1510.06348, ADS entry Cited by: Appendix A.
- Tidal Disruption Events from Stripped Stars. ApJ 973 (1), pp. L9. External Links: Document, 2406.04455, ADS entry Cited by: §5.2.
- The USNO-B Catalog. AJ 125 (2), pp. 984–993. External Links: Document, astro-ph/0210694, ADS entry Cited by: §3.3.
@@
- In-flight calibration of the Swift XRT Point Spread Function. In UV, X-Ray, and Gamma-Ray Space Instrumentation for Astronomy XIV, O. H. W. Siegmund (Ed.), Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 5898, pp. 360–368. External Links: Document, ADS entry Cited by: §B.1.3.
- ZTF discovery of a re-brightening or second transient event towards TDE AT 2021mhg / ZTF21abaxaqq. Transient Name Server AstroNote 138, pp. 1. External Links: ADS entry Cited by: footnote 16.
- XMM-Newton EPIC Background Modeling for Extended Sources. ApJ 629 (1), pp. 172–191. External Links: Document, astro-ph/0504362, ADS entry Cited by: §B.3.1.
- AT 2022upj / ZTF22abegjtx Discovery of extreme coronal lines in nuclear transient. Transient Name Server AstroNote 236, pp. 1. External Links: ADS entry Cited by: §5.4.
- Mapping the Inner 0.1 pc of a Supermassive Black Hole Environment with the Tidal Disruption Event and Extreme Coronal-line Emitter AT 2022upj. ApJ 977 (2), pp. 258. External Links: Document, 2406.11972, ADS entry Cited by: §5.4.
- The tidal disruption event AT2017eqx: spectroscopic evolution from hydrogen rich to poor suggests an atmosphere and outflow. MNRAS 488 (2), pp. 1878–1893. External Links: Document, 1904.10571, ADS entry Cited by: Appendix A, 3rd item, 4th item.
- An outflow powers the optical rise of the nearby, fast-evolving tidal disruption event AT2019qiz. MNRAS 499 (1), pp. 482–504. External Links: Document, 2006.02454, ADS entry Cited by: §1.
- Systematic light-curve modelling of TDEs: statistical differences between the spectroscopic classes. MNRAS 515 (4), pp. 5604–5616. External Links: Document, 2201.02649, ADS entry Cited by: §5.2.
- Secondary standard stars for absolute spectrophotometry.. ApJ 266, pp. 713–717. External Links: Document, ADS entry Cited by: §3.1.
- The nuclear transient AT 2017gge: a tidal disruption event in a dusty and gas-rich environment and the awakening of a dormant SMBH. MNRAS 517 (1), pp. 76–98. External Links: Document, 2206.00049, ADS entry Cited by: Appendix A, 3rd item, §5.4.
- Evolution of Close Binaries. V. The Evolution of Massive Binaries and the Formation of the Wolf-Rayet Stars. Acta Astron. 17, pp. 355. External Links: ADS entry Cited by: §5.2.
- Unveiling the Nature of the Unidentified Gamma-Ray Sources. IV. The Swift Catalog of Potential X-Ray Counterparts. ApJS 209 (1), pp. 9. External Links: Document, 1404.4631, ADS entry Cited by: §B.1.1.
- A Luminous Dust-obscured Tidal Disruption Event Candidate in a Star-forming Galaxy at 42 Mpc. ApJ 948 (1), pp. L5. External Links: Document, 2303.02710, ADS entry Cited by: §6.
- Optical/UV-to-X-Ray Echoes from the Tidal Disruption Flare ASASSN-14li. ApJ 837 (2), pp. L30. External Links: Document, 1703.07024, ADS entry Cited by: Appendix A.
- The Birth of a Relativistic Jet Following the Disruption of a Star by a Cosmological Black Hole. Nature Astronomy 7, pp. 88–104. External Links: Document, 2211.16537, ADS entry Cited by: Appendix A.
- Swift XRT and UVOT deep observations of the high-energy peaked BL Lacertae object PKS 0548-322 close to its brightest state. A&A 462 (3), pp. 889–893. External Links: Document, astro-ph/0610416, ADS entry Cited by: §B.1.3.
- Binary Disruption by Massive Black Holes in Globular Clusters. ApJ 626 (2), pp. 849–852. External Links: Document, astro-ph/0501326, ADS entry Cited by: §4.3.
- SGLF Transient Classification Report for 2023-11-22. Transient Name Server Classification Report 2023-3050, pp. 1. External Links: ADS entry Cited by: Appendix A.
- ASAS-SN Discovery of Luminous Transient in SDSS J110840.11+340552.2. The Astronomer’s Telegram 5831, pp. 1. External Links: ADS entry Cited by: Appendix A.
- The Swift serendipitous survey in deep XRT GRB fields (SwiftFT). I. The X-ray catalog and number counts. A&A 528, pp. A122. External Links: Document, 1102.4270, ADS entry Cited by: §B.1.1.
- Exploring the Optical Transient Sky with the Palomar Transient Factory. PASP 121 (886), pp. 1334. External Links: Document, 0906.5355, ADS entry Cited by: §1.
- Tidal disruption of stars by black holes of 106-108 solar masses in nearby galaxies. Nature 333 (6173), pp. 523–528. External Links: Document, ADS entry Cited by: §1.
- Swift observation of LSQ12dyw. The Astronomer’s Telegram 4346, pp. 1. External Links: ADS entry Cited by: Appendix A.
- The Swift Ultra-Violet/Optical Telescope. Space Sci. Rev. 120 (3-4), pp. 95–142. External Links: Document, astro-ph/0507413, ADS entry Cited by: §1, §3.3.
- XMMSL2 J144605.0+685735: a slow tidal disruption event. A&A 630, pp. A98. External Links: Document, 1908.01065, ADS entry Cited by: §1.
- X-Ray Properties of TDEs. Space Sci. Rev. 216 (5), pp. 85. External Links: Document, ADS entry Cited by: §1.
- First tidal disruption events discovered by SRG/eROSITA: X-ray/optical properties and X-ray luminosity function at z ¡ 0.6. MNRAS 508 (3), pp. 3820–3847. External Links: Document, 2108.02449, ADS entry Cited by: §1, §6.
- Studies in Astronomical Time Series Analysis. VI. Bayesian Block Representations. ApJ 764 (2), pp. 167. External Links: Document, 1207.5578, ADS entry Cited by: §4.1.
- Studies in Astronomical Time Series Analysis. V. Bayesian Blocks, a New Method to Analyze Structure in Photon Counting Data. ApJ 504 (1), pp. 405–418. External Links: Document, astro-ph/9711233, ADS entry Cited by: §4.1.
- Measuring Reddening with Sloan Digital Sky Survey Stellar Spectra and Recalibrating SFD. ApJ 737 (2), pp. 103. External Links: Document, 1012.4804, ADS entry Cited by: §3.3.
- The Man behind the Curtain: X-Rays Drive the UV through NIR Variability in the 2013 Active Galactic Nucleus Outburst in NGC 2617. ApJ 788 (1), pp. 48. External Links: Document, 1310.2241, ADS entry Cited by: §1.
- Release of the ATLAS Forced Photometry server for public use. Transient Name Server AstroNote 7, pp. 1–7. External Links: ADS entry Cited by: §1.
- Delayed appearance and evolution of coronal lines in the TDE AT2019qiz. MNRAS 525 (1), pp. 1568–1587. External Links: Document, 2307.13674, ADS entry Cited by: §5.4.
- PESSTO spectroscopic classification of optical transients. The Astronomer’s Telegram 4299, pp. 1. External Links: ADS entry Cited by: Appendix A.
- Design and Operation of the ATLAS Transient Science Server. PASP 132 (1014), pp. 085002. External Links: Document, 2003.09052, ADS entry Cited by: §1.
- VLASS Tidal Disruption Events with Optical Flares. I. The Sample and a Comparison to Optically Selected TDEs. ApJ 982 (2), pp. 163. External Links: Document, 2310.03791, ADS entry Cited by: §2.1.
- The First Systematically Identified Repeating Partial Tidal Disruption Event. ApJ 985 (2), pp. 175. External Links: Document, 2310.03782, ADS entry Cited by: §4.3, §4.3, footnote 16.
- A tidal disruption event coincident with a high-energy neutrino. Nature Astronomy 5, pp. 510–518. External Links: Document, 2005.05340, ADS entry Cited by: Appendix A, §5.4.
- Lincoln Near-Earth Asteroid Program (LINEAR). Icarus 148 (1), pp. 21–28. External Links: Document, ADS entry Cited by: §1.
- Rates of stellar tidal disruption as probes of the supermassive black hole mass function. MNRAS 455 (1), pp. 859–883. External Links: Document, 1410.7772, ADS entry Cited by: §1.
- Scary Barbie: An Extremely Energetic, Long-duration Tidal Disruption Event Candidate without a Detected Host Galaxy at z = 0.995. ApJ 948 (2), pp. L19. External Links: Document, 2302.10932, ADS entry Cited by: Table 1.